# Circuits and Backdoors: Five Shades of the SETH

### v1

@@ Abstract
The SETH is a hypothesis of fundamental importance to (fine-grained) parameterized complexity theory and many important tight lower bounds are based on it. This situation is somewhat problematic, because the validity of the SETH is not universally believed and because in some senses the SETH seems to be “too strong” a hypothesis for the considered lower bounds. Motivated by this, we consider a number of reasonable weakenings of the SETH that render it more plausible, with sources ranging from circuit complexity, to backdoors for SAT-solving, to graph width parameters, to weighted satisfiability problems. Despite the diversity of the different formulations, we are able to uncover several non-obvious connections using tools from classical complexity theory. This leads us to a hierarchy of five main equivalence classes of hypotheses, with some of the highlights being the following:
  1. 1.
    We show that beating brute force search for SAT parameterized by a modulator to a graph of bounded pathwidth, or bounded treewidth, or logarithmic tree-depth, is actually the same question, despite the structural differences between these parameters. In fact, we show that beating brute force for any of these parameters is possible if and only if one can beat brute force for circuits of depth $\varepsilon n$ (a weaker version of the so-called NC-SETH). We also show that all these questions are equivalent to the existence of an $n^{(1-\varepsilon)k}$ algorithm for the weighted satisfiability problem associated with W[SAT]. A main tool we use here is Barrington’s theorem.
  2. 2.
    On the other hand, modulators to logarithmic pathwidth seem to belong to a different class. We show that beating brute force for this parameter is equivalent to beating brute force for SAT when given a strong 2-SAT backdoor. To prove this we rely on ideas from the classical Immerman–Szelepcsényi theorem.
  3. 3.
    Identifying an even harder class of problems, we show that beating brute force for SAT given a strong Horn backdoor is equivalent to the SETH for general circuits (without depth restrictions) and to beating brute force for the weighted satisfiability problem of the class W[P].
  4. 4.
    Going in the other direction, we revisit recent work of Esmer, Focke, Marx, and Rzazewski (ICALP ’24), and uncover some nuances in their lower bounds. In particular, we show that beating brute force for SAT parameterized by a modulator to constant tree-depth is equivalent to falsifying the standard form of the SETH. This allows us to show that their lower bound for $q$-Coloring (for parameter $(\sigma,\delta)$-hub) is actually not just an implication, but a reformulation of the SETH. On the other hand, we show that their lower bound for Max-Cut for the same parameter is equivalent to the maximization version of the SETH, which at the moment appears to be a weaker hypothesis.
For each of these classes we give at least one (non-SAT) problem which is equivalent to the class as a characteristic example application. Our work is a step towards a more nuanced meta-complexity theory where we advance from proving one-way implications of the SETH, to establishing a classification of interesting questions into equivalence classes. A somewhat unusual aspect is that our classification cuts across traditional time complexity classes, as we are often able to establish equivalences between lower bounds for FPT problems (of the form $(c-\varepsilon)^{k}n^{O(1)}$) and non-FPT problems (of the form $n^{(1-\varepsilon)k}$).
@@ 1.1 Background and Motivation
The Strong Exponential Time Hypothesis (SETH [25]) states that the satisfiability of CNF formulas cannot be decided significantly faster than by brute force, that is, in time $(2-\varepsilon)^{n}$, assuming the formulas have large enough arity [^1] To ease presentation, in this section we use informal statements such as this one to describe all complexity hypotheses. In case of doubt, we refer the reader to Section A.1 where precise definitions are given.. Since its introduction, the SETH has played a major role as an indispensable tool in the development of fine-grained complexity theory. In this paper we focus on its influence on fine-grained parameterized complexity.
@@
The line of work summarized above can be fairly considered quite successful: for many of the most famous problems, not only do we know how to solve them in time $c^{k}n^{O(1)}$ for some constant $c$ and parameter $k$, but we know the precise value of the optimal constant $c$, assuming the SETH. This is great, if one believes the SETH, but it is important to keep in mind that the SETH is not universally believed and indeed, there seem to be serious reasons to doubt its validity, much more than that of more standard assumptions (we refer the reader to the argumentation given by Williams [43] who, somewhat provocatively, assigns the SETH a rather low probability of being true; see also [44]). Although SETH-based lower bounds are still useful (as pointed out in [33] they show that before improving, for instance, the 3-Coloring algorithm we mentioned, we should first try to disprove the SETH), it is somewhat problematic if SETH-based arguments are the only evidence of hardness we have for the most important problems.
This line of reasoning has in fact already been well-developed in the (non-parameterized) fine-grained complexity community. In particular, there are several works which reprove (and improve) lower bounds based on the SETH using a more plausible assumption, such as the NC-SETH, which states the same lower bound but for circuits of poly-logarithmic depth, rather than CNF formulas. We refer the reader to the work of Abboud, Hansen, Vassilevska Williams, and Williams [3] for a more extensive discussion of why the NC-SETH is a preferable hypothesis, or equivalently why falsifying the NC-SETH would have far more surprising consequences than disproving the SETH. The advantages of the NC-SETH over the SETH are taken into account in several other works in this area, for example [1, 9, 41]. In contrast, in parameterized complexity the SETH is often referred to as an “established” hypothesis, though in fairness this is mostly done in comparison to (even more) wide open hypotheses, such as the Set Cover Conjecture [15]. Given the importance of such lower bounds for the wider field, it would be useful if some more convincing evidence could be provided.
@@
The discussion above explains that the SETH is not an ideal starting point for proving lower bounds, as its validity is still subject to some doubt. There is, however, another, more counter-intuitive reason to argue that the SETH is not ideal, and that is that the SETH is “too good” in comparison with the lower bounds we are interested in. This is a line of reasoning recently explored by Esmer, Focke, Marx, and Rzazewski [15], who observed that for many of the problems considered in [33] (including $q$-Coloring, Independent Set, and Max-Cut), a more refined reduction shows that the SETH implies a much stronger lower bound: Not only are current algorithms optimal for treewidth, but they are still optimal for a novel, extremely restricted parameter, which they call $(\sigma,\delta)$-hub. A $(\sigma,\delta)$-hub, (for $\sigma,\delta$ to be considered absolute constants) is a set of vertices (a modulator) whose removal leaves a collection of bounded-size (size $\sigma$) components with bounded degree (degree $\delta$) in the modulator. The class of graphs with a $(\sigma,\delta)$-hub of size $k$ is a tiny fraction of the class of graphs of treewidth $k$, so this is a strong indication that the lower bounds for the more interesting parameters should be achievable by a much weaker assumption.
@@
Coming back to our setting, the above mean that beyond the practical reason we are not satisfied by a lower bound implied by the SETH (we are not sufficiently confident that the SETH is true), there is a philosophical reason not to be happy, which is that it is natural to believe that most of the problems considered are in fact considerably harder than the SETH. It is therefore a natural complexity-theoretic goal to attempt to classify them in a hierarchy of increasing hardness, not just for the added confidence we would gain in the intractability of questions that end up being higher in the hierarchy, but for the extra insight that we are likely to gain from this classification.
@@ 1.2 Summary of Results
One way to explain the main direction of our results is to come back to the recent work of Esmer, Focke, Marx, and Rzazewski [15] we mentioned. We share their initial motivation, which is the observation that the SETH does not just imply tight treewidth-based lower bounds, such as those of [33], but in fact much more. However, where their next step is to say “Let us then try to extract the strongest implication of the SETH”, we say “Let us then explore how we can replace the SETH by a more solid assumption, while still extracting interesting lower bounds”. We will therefore study numerous variants of the SETH, coming from several different areas, including: (i) circuit versions, similar to the NC-SETH (ii) backdoor versions, such as “If we have a SAT formula with a 2-SAT backdoor, can we avoid brute-forcing the backdoor”? (a backdoor is a set of variables such that setting them simplifies the formula) (iii) modulator to graph width versions, such as “If we have a SAT formula with a modulator to a formula of constant treewidth, can we avoid brute-forcing the modulator?” (iv) weighted versions where the question is whether a satisfying assignment of weight $k$ exists (v) maximization versions where the objective is to maximize the number of satisfied clauses.
At this point it is worth pausing and trying to reassure a worried reader. It can be said that one of the weak points of fine-grained complexity theory is the proliferation of reasonable-sounding complexity hypotheses, with no obvious connection to each other. It is natural to think that our decision to study a relatively large and diverse collection of variants of the SETH can lead to nothing more than a chaotic mess, worsening this problem. Thankfully, the main take-home message of our results is that, even though we consider quite a few variations, we are able to identify only five equivalence classes (one of which contains the SETH itself), which align neatly with classes from classical complexity theory. In a sense, the conclusion we draw is optimistic as our results indicate that there is a great deal of complexity-theoretic structure waiting to be discovered here.
Let us then summarize our results, referring also to Figure 1. We identify the following equivalence classes of complexity hypotheses:
@@
    Coming back to the NC-SETH, we see that even though the hypothesis is stated for circuits of poly-logarithmic depth in [3], in practice it is often understood to apply to circuits of depth $n^{o(1)}$ (e.g. [9]). We take this reasoning to its logical conclusion and formulate this hypothesis for circuits of depth $\Omega(n)$, or more precisely for depth at most $\varepsilon n$ (we call this LD-C-SETH, for Linear-Depth Circuits). We discover that this hypothesis is equivalent to stating that brute force is unavoidable for CNF formulas supplied with a modulator to one of three standard widths: (i) constant pathwidth (ii) constant treewidth or (iii) logarithmic tree-depth. We also show that disproving the LD-C-SETH is equivalent to avoiding brute force for the W[SAT]-complete problem of finding a weight-$k$ satisfying assignment to a given (general, non-CNF) formula. This class corresponds to the middle box of Figure 1.
@@
    Going further, we consider the SETH for circuits without any restriction (Circuit-SETH), an assumption which seems much more plausible than the SETH. We show that avoiding brute force in this case can be done if and only if we can avoid brute force in one of two problems (i) deciding satisfiability for a CNF supplied with a Horn backdoor (ii) deciding weight-$k$ satisfiability for an arbitrary circuit, a problem complete for W[P]. This is the top-most class of Figure 1.
@@
    Coming back to the LD-C-SETH, we consider the graph width parameters which are closest the ones already considered. For modulators to logarithmic pathwidth (which would be the logical next step after logarithmic tree-depth) we show that avoiding brute-force enumeration is possible if and only if we can avoid it for 2-SAT backdoors.
@@
    In the other direction, the logical next step below modulators to constant pathwidth is modulators to constant tree-depth. Here we show that the hypothesis that brute-force algorithms are optimal is in fact equivalent to the standard SETH. The parameter $(\sigma,\delta)$-hub introduced by [15] is therefore included here, as it is a restriction of bounded tree-depth modulators. This is the bottom class of Figure 1. In a similar spirit we show that for Max-SAT, the standard version of the hypothesis (for the number of variables) and the version for the size of a $(\sigma,\delta)$-hub are equivalent (this is the second class from the bottom in Figure 1).
[…]
Figure 1: Summary of results. We identify five main equivalence classes, with arrows indicating implication, that is, pointing to weaker hypotheses. From bottom to top the five classes are problems equivalent to the SETH; the SETH for Max-SAT; the SETH for circuits of linear depth; the SETH for 2-SAT backdoors; and the SETH for arbitrary circuits. The figure also shows some connections with related hypotheses. Full definitions are given in Section A.1.
We have therefore identified five natural equivalence classes of hypotheses. In order to make sense of these results, let us attempt to draw an informal parallel with classical complexity theory. All the hypotheses we consider concern problems where the obvious algorithm is a brute-force “guess-and-check”; what differs is the difficulty of the verification procedure. If the verification procedure is P-complete (for example Horn-SAT or evaluating an arbitrary circuit) we end up in the top-most class of Figure 1. Similarly, if it is NL-complete (equivalent to a reachability problem) we end up in the class right below, while the class of the LD-C-SETH seems to correspond to NC1. Finally, the class immediately above the SETH seems somewhat similar to TC0, since in Max-SAT what we need to do if given an assignment is count the number of satisfied clauses and compare this number with a target – this corresponds to the functioning of a threshold gate. Armed with this intuition, we can identify some further complexity questions as characteristic examples of each class.
@@
The initial motivation we gave for this investigation was that we would like to increase our confidence in SETH-based lower bounds for problems parameterized by treewidth. One implication of our results is that we have managed to do so for many problems. Indeed, we show that improving upon the trivial algorithm for $q$-Coloring for parameters modulator size to constant pathwidth, constant treewidth, or logarithmic tree-depth is equivalent to the LD-C-SETH. This implies that improving upon the standard algorithm for pathwidth or treewidth is at least as hard as falsifying the LD-C-SETH (and hence the NC-SETH). As explained, this seems significantly more solid than a result based on the SETH. Furthermore, recent work by Lampis [31] showed that a number of fine-grained lower bounds for problems parameterized by pathwidth (including Independent Set, Dominating Set, Set Cover and others) are equivalent to the pw-SETH, the hypothesis that SAT cannot be solved in time $(2-\varepsilon)^{\textrm{pw}}|\phi|^{O(1)}$, where pw is the input formula’s primal pathwidth. Since the pw-SETH is implied by the same hypothesis for parameter modulator to logarithmic pathwidth, all the hypotheses we consider here except those of the top-most class of Figure 1 also imply the lower bounds of [31]. We have thus succeeded in giving further evidence that known lower bounds for pathwidth and treewidth hold. For instance, even though it was previously known that no $(3-\varepsilon)^{\textrm{pw}}n^{O(1)}$ algorithm is possible for Dominating Set under the SETH, our results (together with those of [31]) establish this under the NC-SETH.
Beyond this, we give a characteristic example problem for each class, based on the intuition linking the classes of Figure 1 with traditional complexity classes. In particular, we show that the Circuit-SETH is equivalent to beating brute-force for the problem of deleting $k$ vertices from a graph, so that it becomes $r$-degenerate, for any fixed $r\geq 2$; this problem is known to be P-complete. Similarly, we give a reachability-type problem ($k$-Neighborhood-Cut) that is equivalent to the 2SatB-SETH, which is natural, since reachability-type problems are usually NL-complete.

### v2

@@ Abstract
The Strong Exponential Time Hypothesis (SETH) is a standard assumption in (fine-grained) parameterized complexity and many tight lower bounds are based on it. We consider a number of reasonable weakenings of the SETH, with sources from (i) circuit complexity (ii) backdoors for SAT-solving (iii) graph width parameters and (iv) weighted satisfiability problems. Our goal is to arrive at formulations which are simultaneously more plausible as hypotheses, but also capture interesting and robust notions of complexity. Using several tools from classical complexity theory we are able to consolidate these numerous hypotheses into a hierarchy of five main equivalence classes of increasing solidity. This framework serves as a step towards structurally classifying a variety of SETH-based lower bounds into intermediate equivalence classes.
To illustrate the applicability of our framework, for each of our classes we give at least one (non-SAT) problem which is equivalent to the class as a characteristic example application. As our main showcase, we consider a natural parameterization of Independent Set by vertex deletion distance from several standard graph classes. Our framework allows us to make connections that would have previously been hard to see: for instance, we consider the question of whether (Weighted) Independent Set can be solved faster than $2^{k}n^{O(1)}$ on graphs which are $k$ vertices away from a class $\mathcal{C}$ where the problem is tractable, such as interval graphs or cographs. We provide precise characterizations of the difficulty of breaking such bounds, in particular proving that obtaining $(2-\varepsilon)^{k}n^{O(1)}$ time algorithms for Cograph+kv+kv or Block+kv+kv graphs is equivalent to obtaining a fast satisfiability algorithm for circuits of depth $\varepsilon n$; while solving the weighted version for Interval+kv+kv graphs is equivalent to the (seemingly) harder problem of obtaining a fast satisfiability algorithm for SAT parameterized by a 2-SAT backdoor.
@@ 1 Introduction
The objective of this paper is to study the structural relations between fine-grained questions in parameterized complexity theory. In order to avoid becoming too abstract and philosophical, let us begin with a concrete example that illustrates what we mean by that.
Consider the Independent Set problem and fix a class of graphs $\mathcal{C}$ on which Independent Set is in P and which satisfies the very mild condition of being hereditary (closed under vertex deletion). How hard is it to solve Independent Set on $\mathcal{C}+kv$, that is, the class of graphs which are $k$ vertex deletions away from being in $\mathcal{C}$? Assuming the $k$ vertices are marked in the input, the obvious time upper bound is $2^{k}n^{O(1)}$: guess which of the $k$ vertices to take in the solution, simplify the instance, and solve the rest in polynomial time. Notice that this is a very standard “distance to triviality” parameterization, of which numerous are considered in the parameterized complexity literature, and that our algorithmic argument is based on a trivial branching procedure.
Given the above, a very natural question to ask is for which classes $\mathcal{C}$ one could hope to do better than $2^{k}n^{O(1)}$. Let us cherry-pick a few example classes which satisfy our criteria, such as interval graphs, cluster graphs, cographs, and block graphs [^1] The precise definitions of these classes are recalled in Section 8, but are not necessary at this point.. Since cluster graphs are a subset of the other classes, one should clearly start there. However, a recent result of Esmer, Focke, Marx, and Rzazewski [19] shows that even for cluster graphs, we cannot achieve $(2-\varepsilon)^{k}n^{O(1)}$, unless the SETH is false [^2] As we explain later this is not explicitly stated in [19] but follows from their construction. We give details in Section 8.! It therefore appears that despite the fact that, say, interval graphs are a much wider class than cluster graphs, the complexity of all these problems (and many other similar cases, say $\mathcal{C}$ being chordal, claw-free, $P_{5}$-free, etc.) simply collapse to one identical case. Should we view this as cause for celebration (because we have completely resolved all these problems)? Or should we view this as cause for concern (are we missing something)?
Before we answer this, let us take a step back into classical complexity theory. Consider the following questions: we are given a CNF formula $\phi$ and are asked (i) is $\phi$ satisfiable? (ii) does $\phi$ have a unique satisfying assignment? (iii) is the majority of assignments satisfying for $\phi$? (iv) can we satisfy $\phi$ even if we alternate picking values for variables with an adversary who is trying to falsify it? A knowledgeable reader can easily recognize that the complexity of these questions is strongly believed to be very different, as they correspond to complete problems for the classes NP, $\Sigma_{2}^{p}$, PP, and PSPACE respectively. Nevertheless, a reader less familiar with complexity theory may see a similarity between this situation and the questions of the previous paragraph. Indeed, all the questions we asked about $\phi$ can be solved in exponential time (and polynomial space), while exponential time is necessary even for the easiest one (under standard hypotheses). We observe then that a large effort of classical complexity theory has been devoted into understanding the relative structural complexities of various problems, even (and perhaps especially) in cases where the problems are believed to have the same time and space complexity. This effort has been quite impactful as the different complexity nuances captured by all the intermediate classes have often turned out to be important in other aspects (for instance, kernelization lower bounds typically rely on the assumption that the polynomial hierarchy does not collapse, which would not make sense as a statement if we only cared about deterministic time complexity).
Given this history, the thesis of this paper is that if, in the context of fine-grained parameterized complexity, we settle just for SETH-based lower bounds, then we are indeed missing something. Going back to our example, we would like to replicate this success of classical complexity theory in our setting and obtain a more nuanced understanding of the complexity of our questions. In the same way that traditional complexity theory allows us to go from saying “this problem is likely intractable (not in P)” to saying “this problem is exactly as hard as $\Sigma_{2}^{p}$ or PSPACE”, we would like to be able to go from saying “brute-force is inevitable for Independent Set on the class $\mathcal{C}+kv$ (under SETH)” to “avoiding brute force for Independent Set on $\mathcal{C}+kv$ is exactly as hard as this class of problems”. The obstacle is of course that, at the moment, fine-grained parameterized theory is missing the framework and the vocabulary necessary to do this. In other words, we do not yet have a classification effort which would allow us to take questions which are all “SETH-hard” and order them into equivalence classes, according to their relative difficulty.
The goal of this paper is exactly to explore this new direction, start building the basic blocks of such a framework, and show that this is a fruitful and promising approach. We will do this by identifying five natural equivalence classes (Figure˜1), one of which contains the SETH, and giving natural complete problems for each as well as an informal characterization connecting them to classical complexity classes. To showcase the applicability of this framework we will also apply it to the Independent Set problem discussed above for the four classes we cherry-picked (Figure˜2). The framework will allow us to compare the relative complexities of cases which do not have an obvious graph-theoretic relation (for example cographs and interval graphs are incomparable classes), but also relate them with questions that initially seem completely unrelated, stemming from circuit complexity or backdoors for SAT-solving. We see this investigation as a first step that will allow us to more carefully classify known tight lower bounds from the literature into equivalence classes, with the added benefit of identifying the lower bounds which should be considered most plausible (because it must be taken into account that the SETH is not a universally believed assumption).
@@ 1.1 Background and Motivation
Having summarized the general direction of this paper, let us give a more detailed review of the relevant background. Recall that the Strong Exponential Time Hypothesis (SETH [30]) states that the satisfiability of CNF formulas cannot be decided significantly faster than by brute force, that is, in time $(2-\varepsilon)^{n}$, assuming the formulas have large enough arity [^3] To ease presentation, in this section we use informal statements such as this one to describe all complexity hypotheses. In case of doubt, we refer the reader to Appendix A, where precise definitions are given.. Since its introduction, the SETH has been an indispensable tool in the development of fine-grained complexity theory. In this paper we focus on its influence on fine-grained parameterized complexity.
@@
One of the motivations of this paper is to question the wisdom of basing all these numerous lower bound results on one-way implications of the SETH. Even though the SETH is mostly considered a reasonable hypothesis, it is important to keep in mind that it is not universally believed. Indeed, there seem to be serious reasons to doubt its validity, much more than that of more standard assumptions (we refer the reader to the argumentation given by Williams [49] who, somewhat provocatively, assigns the SETH only a one-in-four probability of being true; see also [50]). In other words, even if one believes that the SETH is probably true, because its status is sufficiently open one should not necessarily consider a SETH-based lower bound as the final word on a problem’s status.
This line of reasoning has already been well-developed in the (non-parameterized) fine-grained complexity community. In particular, there are several works which reprove (and improve) lower bounds based on the SETH using a more plausible assumption, such as the NC-SETH, which states the same lower bound but for circuits of poly-logarithmic depth, rather than CNF formulas. We refer the reader to the work of Abboud, Hansen, Vassilevska Williams, and Williams [3] for a more extensive discussion of why the NC-SETH is a preferable hypothesis, or equivalently why falsifying the NC-SETH would have far more surprising consequences than disproving the SETH; see also [1, 10, 47]. In the parameterized complexity setting, similar work was only recently initiated in [36] (more on this below).
@@
The discussion above explains that the SETH is not an ideal starting point for proving lower bounds, as its validity is still subject to some doubt. There is, however, another, more counter-intuitive reason to argue that the SETH is not ideal, and that is that the SETH is “too good” in comparison with the lower bounds we are interested in. This is a line of reasoning taken by several previous works [26, 32]. In particular, Jaffke and Jansen [32] showed that the correct base for $q$-Coloring is exactly $q$ not only when we parameterize by pathwidth or treewidth, but even when we parameterize by vertex deletion distance to path-forest (a significantly more restrictive parameter). More recently, this approach was further explored by Esmer, Focke, Marx, and Rzazewski [19], who observed that for many of the problems considered in [38] (including $q$-Coloring, Independent Set, and Max-Cut), a more refined reduction shows that the SETH implies a much stronger lower bound: Not only are current algorithms optimal for treewidth, but they are still optimal for a novel, extremely restricted parameter, which they call $(\sigma,\delta)$-hub. A $(\sigma,\delta)$-hub, (for $\sigma,\delta$ to be considered absolute constants) is a set of vertices (a modulator) whose removal leaves a collection of bounded-size (size $\sigma$) components with bounded degree (degree $\delta$) in the modulator. The class of graphs with a $(\sigma,\delta)$-hub of size $k$ is a tiny fraction of the class of graphs of treewidth $k$, so this is a strong indication that the lower bounds for the more interesting parameters should be achievable by a much weaker assumption.
Another way to see the above is that the SETH is talking about the hardness of a different class of problems than the width-based lower bounds that it implies. Indeed, observe that we have crucially moved from problems where the “obvious” algorithm is dynamic programming (parameterized by treewidth or other widths) to problems where the “obvious” algorithm is branching (for example, decide the color of the vertices of the modulator). This leads us to believe that there are several “complexity classes” lying between the SETH and the treewidth-based lower bounds discussed above and waiting to be discovered.
@@
Much of the state of the art in fine-grained parameterized complexity can then be thought of as being close to the superficial understanding we sketched above: we show that our problems are harder than the SETH, and consider the case closed. The thesis of this paper is that there is something to be gained by more carefully comparing the relative complexities of our problems, and for this we need to organize them into equivalence classes. Interestingly, we will show that the equivalence classes we end up with can sometimes be orthogonal to traditional complexity classes.
[…]
Figure 1: Summary of results. We identify five main equivalence classes, with arrows indicating implication, that is, pointing to weaker hypotheses. From bottom to top the five classes are problems equivalent to the SETH; the SETH for Max-SAT; the SETH for circuits of linear depth; the SETH for 2-SAT backdoors; and the SETH for arbitrary circuits. The figure also shows some connections with related hypotheses. Full definitions are given in Appendix˜A.
@@ 1.2 Summary of Results
One way to explain the main direction of our results is to come back to the recent work of Esmer, Focke, Marx, and Rzazewski [19] we mentioned. We share their initial motivation, which is the observation that the SETH does not just imply tight treewidth-based lower bounds, such as those of [38], but in fact much more. However, where their next step is to say “Let us then try to extract the strongest implication of the SETH”, we say “Let us then explore how we can replace the SETH by a different (more solid) hypothesis, and try to understand what type of question we capture”. We will therefore study numerous variants of the SETH, coming from several different areas, including: (i) circuit versions, similar to the NC-SETH (ii) backdoor versions, such as “If we have a SAT formula with a 2-SAT backdoor, can we avoid brute-forcing the backdoor”? (a backdoor is a set of variables such that setting them simplifies the formula) (iii) modulator to graph width versions, such as “If we have a SAT formula with a modulator to a formula of constant treewidth, can we avoid brute-forcing the modulator?” (iv) weighted versions where the question is whether a satisfying assignment of weight $k$ exists (v) maximization versions where the objective is to maximize the number of satisfied clauses.
At this point it is worth pausing and trying to reassure a worried reader. One of the weak points of fine-grained complexity theory is the proliferation of reasonable-sounding complexity hypotheses, with no obvious connection to each other. It is natural to think that our decision to study a relatively large and diverse collection of variants of the SETH can only make things worse. Thankfully, the main take-home message of our results is that, even though we consider quite a few variations, we are able to identify only five equivalence classes (one of which contains the SETH itself), which align neatly with classes from classical complexity theory. In a sense, the conclusion we draw is optimistic as our results indicate that there is a great deal of complexity-theoretic structure waiting to be discovered here.
Let us then summarize our results, referring also to Figure˜1. We identify the following equivalence classes of complexity hypotheses:
@@
    Possibly the most interesting of the classes we consider is the middle one of Figure˜1, which is motivated by the NC-SETH. Even though the NC-SETH is stated for circuits of poly-logarithmic depth in [3], in practice it is often understood to apply to circuits of depth $n^{o(1)}$ (e.g. [10]). We take this reasoning to its logical conclusion and formulate this hypothesis for circuits of depth $\Omega(n)$, or more precisely for depth at most $\varepsilon n$ (we call this LD-C-SETH, for Linear-Depth Circuits). We discover that this hypothesis is equivalent to stating that brute force is unavoidable for CNF formulas supplied with a modulator to one of three standard widths: (i) constant pathwidth (ii) constant treewidth or (iii) logarithmic tree-depth. We find it somewhat surprising that these parameters, which are know to strictly generalize each other in the graph-theoretic sense, collapse to the same hypothesis. We also show that disproving the LD-C-SETH is equivalent to avoiding brute force for the W[SAT]-complete problem of finding a weight-$k$ satisfying assignment to a given (general, non-CNF) formula. This class corresponds to the middle box of Figure˜1. An important remark here is that we are able to extend these results to MaxSAT (see Section˜4.4), as this requires a foray into arithmetic circuits and is crucial later on to obtain the precise characterization of the complexity of Independent Set on graphs which are close to being cographs or block graphs. This class therefore seems to be quite robust and contains some very natural equivalent problems unrelated to SAT.
@@
    Going further, we consider the SETH for circuits without any restriction (Circuit-SETH), an assumption which seems much more plausible than the SETH. We show that avoiding brute force in this case can be done if and only if we can avoid brute force in one of two problems (i) deciding satisfiability for a CNF supplied with a Horn backdoor (ii) deciding weight-$k$ satisfiability for an arbitrary circuit, a problem complete for W[P]. This is the top-most class of Figure˜1.
@@
    Coming back to the LD-C-SETH, we consider the graph width parameters which are closest to the ones already considered. For modulators to logarithmic pathwidth (which would be the logical next step after logarithmic tree-depth) we show that avoiding brute-force enumeration is possible if and only if we can avoid it for 2-SAT backdoors. As we explain below, one of our main showcase results is to show that an equivalent characterization of this class can be given by Weighted Independent Set on $\textrm{Interval}+kv$ graphs, so this class also seems quite natural.
@@
    In the other direction, the logical next step below modulators to constant pathwidth is modulators to constant tree-depth. Here we show that the hypothesis that brute-force algorithms are optimal is in fact equivalent to the standard SETH. The parameter $(\sigma,\delta)$-hub introduced by [19] is therefore included here, as it is a restriction of bounded tree-depth modulators. In other words, our main result here is that hypotheses based on any parameter more restricted than “modulator to bounded tree-depth” all collapse to the normal SETH. This is the bottom class of Figure˜1. In a similar spirit we show that for Max-SAT, the standard version of the hypothesis (for the number of variables) and the version for the size of a $(\sigma,\delta)$-hub are equivalent (this is the second class from the bottom in Figure˜1).
Take-home message
We have identified five natural equivalence classes of hypotheses. In order to make sense of these results, let us attempt to draw an informal analogy with classical complexity theory. All the hypotheses we consider concern problems where the obvious algorithm is a brute-force “guess-and-check”; what differs is the difficulty of the verification procedure. If the verification procedure is P-complete (for example Horn-SAT or evaluating an arbitrary circuit) we end up in the top-most class of Figure˜1. Similarly, if it is NL-complete (equivalent to a reachability problem) we end up in the class right below, while the class of the LD-C-SETH seems to correspond to NC1. Finally, the class immediately above the SETH seems somewhat similar to (a restricted version of) TC0, since in Max-SAT what we need to do if given an assignment is count the number of satisfied clauses and compare this number with a target – this corresponds to the functioning of a threshold gate. What we are then essentially doing is reconstructing a parallel of a part of traditional complexity theory around fine-grained variations of the SETH and what remains is to populate these classes with further problems. Armed with this intuition, we can identify some further complexity questions as characteristic examples of each class.
@@
One motivation we gave for this investigation was that we would like to increase our confidence in SETH-based lower bounds for problems parameterized by treewidth. One implication of our results is that we have managed to do so for many problems. Indeed, we show that improving upon the trivial algorithm for $q$-Coloring for parameters modulator size to constant pathwidth, constant treewidth, or logarithmic tree-depth is equivalent to the LD-C-SETH. This implies that improving upon the standard algorithm for pathwidth or treewidth is at least as hard as falsifying the LD-C-SETH (and hence the NC-SETH). As explained, this seems significantly more solid than a result based on the SETH. Furthermore, recent work by Lampis [36] showed that a number of fine-grained lower bounds for problems parameterized by pathwidth (including Independent Set, Dominating Set, Set Cover and others) are equivalent to the pw-SETH, the hypothesis that SAT cannot be solved in time $(2-\varepsilon)^{\textrm{pw}}|\phi|^{O(1)}$, where pw is the input formula’s primal pathwidth. Since the pw-SETH is implied by the same hypothesis for parameter modulator to logarithmic pathwidth, all the hypotheses we consider here except those of the top-most class of Figure˜1 also imply the lower bounds of [36]. We have thus succeeded in giving further evidence that known lower bounds for pathwidth and treewidth hold. For instance, even though it was previously known that no $(3-\varepsilon)^{\textrm{pw}}n^{O(1)}$ algorithm is possible for Dominating Set under the SETH, our results (together with those of [36]) establish this under the NC-SETH.
Beyond this, we give a characteristic example problem for each class, based on the intuition linking the classes of Figure˜1 with traditional complexity classes. In particular, we show that the Circuit-SETH is equivalent to beating brute-force for the problem of deleting $k$ vertices from a graph, so that it becomes $r$-degenerate, for any fixed $r\geq 2$; this problem is known to be P-complete. Similarly, we give a reachability-type problem ($k$-Neighborhood-Cut) that is equivalent to the 2SatB-SETH, which is natural, since reachability-type problems are usually NL-complete.